x^2+14x+40=0 find all solutions
Answer by jim_thompson5910(35256)


x^2+14x+40=0

(x+10)(x+4)=0

x+10 = 0 or x+4=0

x = -10 or x = -4

The two solutions are x = -10 or x = -4